A particle of mass m moves under the influence of gravity along the helix z =kθ, r =constant, where k is a constant and z is vertical. Obtain the Hamiltonian equations of motion.


1
Expert's answer
2021-12-06T09:41:52-0500

"z=k\theta,"

"L=T-U,"

"T=\frac{km\theta^2}2,"

"U=kmg\ddot{\theta},"

"\frac{\partial L}{\partial \theta}-\frac d{dt}\frac{\partial L}{\partial \dot {\theta}}=0,"

"km\theta\dot{\theta}-mg\ddot{\theta}=0."